The FBI wants to determine the effectiveness of their 10 Most Wanted list. To do so, they need to find out the fraction of people who appear on the list that are actually caught.

Step 1 of 2: Suppose a sample of 369 suspected criminals is drawn. Of these people, 118 were captured. Using the data, estimate the proportion of people who were caught after being on the 10 Most Wanted list. Enter your answer as a fraction or a decimal number rounded to three decimal places.

______

Step 2 of 2: Suppose a sample of 369  suspected criminals is drawn. Of these people, 118 were captured. Using the data, construct the 90% confidence interval for the population proportion of people who are captured after appearing on the 10 Most Wanted list. Round your answers to three decimal places.

Lower Endpoint:_____ Upper Endpoint:_____
Answer #1

a) The sample proportion is:

\hat{p}=\frac{x}{n}=\frac{118}{369}={\color{Red} 0.320}

------------------------------------------------------------------------------------------------------

b) For 90% confidence level, critical value of z is z^*=1.645 .

\\Lower \ endpoint=\hat{p}- z^* \sqrt{\frac{\hat{p}(1-\hat{p})}{n}}=0.320-1.645\sqrt{\frac{0.320(1-0.320)}{369}} \\\\=0.320-0.040={\color{Red} 0.280}

\\Upper \ endpoint=\hat{p}+ z^* \sqrt{\frac{\hat{p}(1-\hat{p})}{n}}=0.320+1.645\sqrt{\frac{0.320(1-0.320)}{369}} \\\\=0.320+0.040={\color{Red} 0.360}